OpenStudy (anonymous):

Bettina's age is three times Melvina's. If 20 is added to Melvina's age and 20 is subtracted from Bettina's, their ages will be equal. How old is each now?

OpenStudy (anonymous):

B = 3M--------------(1) And, M+20 = B-20 or, B- M=40 or, 3M -M= 40------ from (1) or, 2M= 40 or, M = 20 And, B= 3M= 3 * 20 = 60